ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: Barbu, Luminita, Constanta, and Gheorghe Morosanu, Iasi, Romania Asymptotic Analysis of the Telegraph Equations The behaviour of the solution of a nonlinear non-local boundary value problem associated to the telegraph equations with small inductance is studied.